OpenStudy (anonymous):

Which ratio is not equivalent to Proportion 24/36 ? A. Proportion B. Proportion C. Function Equations D. Proportion

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Seems like you have a copy paste error. What are the answer choices?

OpenStudy (anonymous):

http://prntscr.com/6qshh8

TheSmartOne (thesmartone):

First lets simplify \(\sf\Large\frac{24}{36}\) That equals to \(\sf \Large\frac{4 \times 6}{6 \times 6}\) So we can divide both sides by 6, and then what will it be simplified to?

OpenStudy (anonymous):

D

TheSmartOne (thesmartone):

Good job :)